OpenStudy (anonymous):

If I have a "Dividing Radical Expressions" question and the denominator has something like 4+4 square roots of 5, would I need to times both of those numbers by the square root of 5?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

Can you be more specific about what you're looking at? I can't quite understand your description.

OpenStudy (anonymous):

multiply top and bottom by\[4-\sqrt{5}\]

OpenStudy (anonymous):

\[(4+\sqrt{5})(4-\sqrt{5})=16-5=11\]

OpenStudy (anonymous):

so my equation is 3/ 4+4 square roots of 5. In order to solve the problem, there can be no radicals on the bottom half. Would I need to times both the 4 and the 4 square roots of 5 by the square root of 5?

OpenStudy (anonymous):

Oh, as Satellite said. Multiply by the conjugate. \[(a + \sqrt{b})(a - \sqrt{b}) = a^2 -b\] So if you have \(a + \sqrt{b}\) you would multiply by \(a - \sqrt{b}\) and vice versa.

OpenStudy (anonymous):

If you did that, you would get: \[3/(4+4\sqrt{5})=3 \sqrt{5}/(20+4\sqrt{5})\] This does not cancel the square roots in the denominator so, instead, multiply the top and bottom by the conjugate of the denominator: \[4-4\sqrt{5}\] This gives: \[3/(4+4\sqrt{5})=3(4-4\sqrt{5})/(4+4\sqrt{5})(4-4\sqrt{5})=(12-12\sqrt{5})/(-64)\] There are no square roots in the denominator of this expression so your problem is solved.

OpenStudy (anonymous):

So what if the equation was that the alone 4 was negative

OpenStudy (anonymous):

Wouldn't matter.

OpenStudy (anonymous):

The part you change the sign on is the part with the radical.

OpenStudy (anonymous):

sorry i did not read carefully. if it was \[4+4\sqrt{5}\] you multiply by \[4-4\sqrt{5}\]

OpenStudy (anonymous):

When I use the conjugated form, do I need to change the negative 3 to a positive 3?

OpenStudy (anonymous):

multiply by the 'conjugate'

OpenStudy (anonymous):

conjugate of \[a+b\] is \[a-b\]

OpenStudy (anonymous):

what if it was negitive a + b? then what would it be?

OpenStudy (anonymous):

With square roots, the conjugate of \[a +\sqrt{b}\] is \[a-\sqrt{b}\]Basically, change the sign on the square root to find the conjugate.

OpenStudy (anonymous):

the point being that \[(a+b)(a-b)=a^2-b^2\] so for example \[(3+\sqrt{2})(3-\sqrt{2})=3^2)-(\sqrt{2})^2=9-2=7\]

OpenStudy (anonymous):

or \[(\sqrt{5}-3)(\sqrt{5}+3)=(\sqrt{5})^2-3^2=5-9=-4\] etc

OpenStudy (anonymous):

Yes, multiplying by a conjugate always cancels the square roots.